The sum to n terms of an arithmetic series is sn=n/2 +(7n+15).determine the 6th tern and how many terms of the series must be added to give a sum of 425
Answer by richwmiller(17219)

Are you sure it is Sn=n/2 +(7n+15) plus and not Sn=n/2*(7n+15) times?